# How LLMs Learn to Reason: A Complex Network Perspective

### v1

@@ Abstract
Training large language models with Reinforcement Learning from Verifiable Rewards (RLVR) exhibits a set of distinctive and puzzling behaviors that remain poorly understood, including a two-stage learning curve, V-shaped response-length trajectories, and a pronounced vulnerability to catastrophic forgetting. In this work, we propose that these seemingly disparate phenomena can be explained using a single unifying theory: the model’s reasoning process maps to the self-organization of a semantic complex network whose topology remains persistently sparse, with the average degree pinned close to two. This topology imposes a fundamental mechanism for forgetting and learning: it first drives the system into a maximally frustrated state where “skill islands” form, slow-learning happens, and forgetting is induced; then it enters a sharp growth phase where the new skills are “bolted on”, driven by phase-transition-like learning at the web’s frontier. Equipped with the theory, we propose Annealed-RLVR, a principled algorithm that introduces an SFT-based “heating” step at the point of maximal frustration to resolve the competitive bottleneck and enhance the reasoning capability of the model. Experiments on a 1.5B-parameter model demonstrate that the approach outperforms standard RLVR on both in-distribution and out-of-distribution benchmarks. By recasting RLVR from black-box optimization into a predictable process of structural self-organization, our work provides a new physical intuition for engineering the emergent reasoning capabilities of future AI systems.
@@ 1 Introduction
Large language models (LLMs) trained with Reinforcement Learning from Verifiable Rewards (RLVR) (Liu et al., 2025; Guo et al., 2025; Team et al., 2025; Yang et al., 2025) are increasingly capable of multi-step, “System-2” reasoning. Yet, this training paradigm consistently exhibits a set of puzzling macroscopic behaviors across diverse architectures and benchmarks. These include (i) rapid initial gains that give way to a prolonged plateau; (ii) a V-shaped trajectory in the length of correct solutions, which first shortens before steadily lengthening (Luo et al., 2025; He et al., 2025; Fatemi et al., 2025); and (iii) a pronounced vulnerability to catastrophic forgetting when interleaved with supervised fine-tuning (SFT) (Li & Hoiem, 2017; Fernando et al., 2025; Ding & Wang, 2025). Although these phenomena are widely observed, the reasons for their co-occurrence remain elusive. A clear understanding of the underlying mechanism would be essential for developing more efficient approaches to training reasoning-capable models.
@@
A promising theoretical framework analyzes these phenomena by modeling the LLM’s reasoning process on an implicit reasoning graph — a network of semantic states and logical transitions(Wang et al., 2024; Dutta et al., 2024; Xiang et al., 2025; Cabannes et al., 2024; Minegishi et al., 2025). However, this perspective faces a critical roadblock: a direct, microscopic analysis is currently intractable. Constructing this entire graph from a model’s high-dimensional latent space remains a formidable open problem, preventing a direct investigation into the structural origins of the observed dynamics.
@@
Here, we argue that the very universality of the RLVR phenomena provides a path forward. In complex systems, behaviors that are independent of microscopic details are often emergent properties, governed by the system’s large-scale organization (Anderson, 1972; Watts & Strogatz, 1998; Barabási & Albert, 1999). This insight motivates an approach rooted in the core logic of the Renormalization Group (RG)(Wilson & Kogut, 1974): instead of getting bogged down by intractable local details, one can leap directly to hypothesizing a core principle about the system’s large-scale structure to explain its macroscopic behavior. Applying this strategy, we shift our focus from the full reasoning graph to its coarse-grained backbone, which emerges by abstracting away fine-grained skills encoded in local network patterns like motifs or community structures. This leads us to our central hypothesis regarding the emergent structure that forms under RLVR training for general-purpose System-2 reasoning:
@@
Figure 1: CoNet Faithfully Reproduces Core LLM Training Dynamics. The minimal CoNet model (a) remarkably reproduces the two core empirical signatures of RLVR training observed in the DeepScaleR-1.5B LLM (b). These signatures are: (i) a two-stage reward dynamic consisting of a fast-learning phase followed by a slow-learning plateau (top panels), and (ii) a non-monotonic, V-shaped evolution of the correct response length (bottom panels). This striking correspondence validates CoNet as a tractable minimal model for our theoretical analysis. The inset in (a) schematically depicts the CoNet abstraction. Lines in (b) are smoothed with a 60-step moving average.
To validate this hypothesis, we combine two approaches: a theoretical study using the Concept Network Model (CoNet) (Cai et al., 2025), a minimal “computational microscope” that reproduces the LLM’s core training dynamics, and numerical experiments on a 1.5B-parameter LLM(Luo et al., 2025). This dual approach reveals a clear structural narrative for the observed puzzles. The learning plateau is shown to be the arduous consolidation of disconnected “skill islands”, a transition that directly explains the V-shaped length trajectory. Furthermore, the resulting sparse, tree-like concept web accounts for the model’s fragility and its vulnerability to catastrophic forgetting with SFT.
This sparse topology imposes a microscopic dynamic of learning and forgetting that is a hallmark of the RLVR-driven self-organization process. We find the degree-2 constraint first forces the system into a maximally frustrated state at the onset of the slow-learning plateau, a phase dominated by frustration-induced forgetting as skills are competitively suppressed. Subsequently, the very same structure enables the growth mechanism that governs the entire slow-learning phase: a steady, sequential process where new skills are individually “bolted on” at the web’s expanding frontier via a phase-transition-like learning dynamic.
Equipped with the sparse-web theory, we propose Annealed-RLVR, a principled algorithm that introduces an SFT-based “heating” step at the point of maximal frustration to resolve the competitive bottleneck and enhance exploration and reasoning efficacy of the subsequent RLVR phase. Experiments on both CoNet and the 1.5B-parameter LLM confirm the efficacy of this strategy: Annealed-RLVR surpasses standard RLVR on both in-distribution and out-of-distribution benchmarks, transforming our physical intuition into a practical tool for engineering emergent reasoning.
@@
Figure 2: Visualizing the Structural Evolution from Skill Islands to the Concept Web. These network snapshots provide a direct, qualitative narrative of the concept web’s formation, corresponding to the different phases of training. (a) In the early fast-learning phase (Step 20), the model discovers a few short, disjointed reasoning paths, representing the first nascent “skill islands”. (b) At the onset of slow learning (Step 50), the system has proliferated into a maximal collection of disconnected islands. (c) Deep in the slow-learning phase (Step 800), these previously separate islands have coalesced into a single, giant connected component, forming the unified and expansive concept web.
@@ 2 The V-Shaped Signature as Evidence for a Sparse Concept Web
The V-shaped evolution of response length is a well-demonstrated, quantifiable hallmark of deep structural reorganization during RLVR training(He et al., 2025; Luo et al., 2025). To anchor our investigation, we first reproduce this signature by applying RLVR to the DeepSeek-R1-Distill-Qwen-1.5B model following the DeepScaleR protocol (Luo et al., 2025). As shown in Fig. 1, our experiment recovers the characteristic V-shaped curve (Fig. 1). We contend that this is not a mere by-product but the primary empirical evidence for our central claim: the emergence of a sparse, unified concept web.
To dissect the mechanism underlying this signature—a task intractable within the LLM itself—we deploy an extended, multi-task variant of the CoNet as a “computational microscope” (See Appendix D for details). This minimal model reduces the learning process to a tractable graph-traversal problem and, remarkably, reproduces the V-shaped curve observed in our LLM experiment (Fig. 1), validating it as a faithful proxy for uncovering the underlying mechanism.
Armed with this confirmation, we now expose the topological origin of the V-shape. The V-shaped curve arises from a distinct two-stage process. During the initial fast-learning phase where response length decreases, the model undergoes local optimization: it rapidly discovers short, efficient solutions to individual problems, giving rise to a proliferation of disconnected “skill islands” (as shown in Fig. 2(a)). This burst is quantified by the number of distinct solution clusters, which peaks right before the onset of slow learning (Fig. 2(b)). Stage one thus constitutes a parallel search for local efficiency that naturally drives down the average response length.
@@
Figure 3: A Sparse Web Structure Necessitates Longer Reasoning Chains. This figure links the emergent sparse topology of the concept web to the observed increase in response length. (b, c) The largest connected component (edges with probability > 0.95) remains sparse (average degree $\approx 2$) even as it grows from step 50 to 800. (a) Consequently, the distribution of solution path lengths shifts decisively to the right during this period, confirming that navigating the sparse backbone requires longer reasoning chains.
@@
The subsequent slow-learning phase, where response length steadily increases, marks a fundamental shift to global integration. The primary task is no longer discovering new islands but weaving them into a single, expansive concept web (see Fig. 2(c)). As the largest connected component relentlessly grows, a critical structural property emerges: the web remains persistently sparse, with its average degree stabilizing near two (as shown in Fig. 3(c)). This provides direct, quantitative validation for our central Hypothesis (Hypothesis 1).
@@
This emergent sparsity is the topological key to the V-shape’s rising slope. A network with such a low average degree is dominated by chain and tree-like structures, inherently lacking the shortcuts and redundant pathways found in denser graphs. Consequently, connecting two previously distant concepts requires traversing long, intermediate paths. The histograms in Fig. 3(a) provide dispositive proof of this effect: As the sparse web expands, the distribution of solution lengths undergoes a decisive and sustained rightward shift, confirming that navigation on this sparse backbone necessitates longer reasoning chains. Thus, the V-shaped curve furnishes a robust macroscopic signature of an underlying topological transition: from parallel, local optimization on disconnected components to serial, global navigation across an expanding, sparse backbone.
@@
Yet can this simple, largely tree-like model capture the full complexity of LLM reasoning? After all, an LLM’s autoregressive memory enables sophisticated behaviors such as self-correction or reflection (e.g., “Wait, let me rethink that…”)—that would manifest as cycles in the reasoning graph. We argue they are fully compatible(Minegishi et al., 2025). These reasoning cycles are not the dense, ubiquitous shortcuts of a highly connected graph. Instead, they are likely the cognitive manifestation of sparse, local loops—occasional, high-cost revisions on a much larger, fundamentally chain-like scaffold. The existence of a limited number of such cycles does not significantly alter the global average degree of the web, which remains close to two. With the sparse web model established, this link between a macroscopic dynamic and a microscopic topology forms the foundation for the analysis that follows.
@@
Figure 4: The Fragility of a Sparse Web: Catastrophic Forgetting and Fast Recovery. This figure demonstrates a key prediction of our sparse web hypothesis: that the concept web is fragile, relying on critical bridge-like connections. (c-e) The microscopic view in CoNet shows how a lightweight supervised fine-tuning (SFT) on a converged web (c) severs these critical bridges, causing the structure to fragment (d). Subsequent RLVR rapidly repairs these links (e). (a, b) This microscopic severing manifests as macroscopic catastrophic forgetting in both CoNet and the 1.5B LLM, where performance collapses upon initiating SFT. The subsequent fast recovery once RLVR resumes highlights the localized nature of the damage, underscoring the web’s structural fragility.

### v2

@@ Abstract
Training large language models with Reinforcement Learning with Verifiable Rewards (RLVR) exhibits a set of distinctive and puzzling behaviors that remain poorly understood, including a two-stage learning curve, a V-shaped response-length trajectory, and a pronounced vulnerability to catastrophic forgetting. In this work, we propose that these behaviors are emergent collective phenomena governed not by neural implementation details, but by the topological evolution of the latent reasoning graph in semantic space. By demonstrating a dynamical isomorphism between a 1.5B-parameter LLM and a minimal Concept Network Model (CoNet), we trace the causal source to the self-organization of a sparse concept web pinned to an average degree of two. This geometric perspective provides a unified physical explanation for the observed anomalies: the V-shaped trajectory tracks the evolution from parallel local skill optimization to global network integration; catastrophic forgetting stems from the topological disconnection of critical “trunk” edges; and policy collapse arises from the accumulation of sequential transitions at the web’s leaf nodes, where broad exploration abruptly freezes into rigid, high-reward trajectories. Identifying a “maximally frustrated state” at the transition between learning stages, we propose Annealed-RLVR, a principled algorithm that injects a targeted SFT “heating” step to resolve this topological bottleneck. Experiments confirm that this theory-driven intervention outperforms standard RLVR on both in-distribution and out-of-distribution benchmarks (including Minerva and AIME). By recasting RLVR from black-box optimization into a predictable process of structural self-organization, our work provides a new physical intuition for engineering the emergent reasoning capabilities of future AI systems.
@@ 1 Introduction
Training large language models with reinforcement learning with verifiable rewards (RLVR) (Liu et al., 2025; Guo et al., 2025; Team et al., 2025; Yang et al., 2025) has revealed four striking and still poorly understood phenomena. On math and coding benchmarks, RLVR runs typically show (i) rapid initial gains followed by a long performance plateau, even as training continues (Luo et al., 2025a; He et al., 2025; Jin et al., 2025b); (ii) a V-shaped trajectory of the length of correct solutions, where answers first become markedly shorter before later growing longer again (Luo et al., 2025a; He et al., 2025; Fatemi et al., 2025; Wang et al., 2025); (iii) pronounced vulnerability to catastrophic forgetting when RLVR-trained policies are interleaved or followed by supervised fine-tuning (SFT), with earlier reasoning skills abruptly degraded (Li & Hoiem, 2017; Luo et al., 2025b; Ding & Wang, 2025); and (iv) a collapse of solution diversity, with the policy converging onto a narrow and brittle solution manifold (He et al., 2025; Cui et al., 2025; Hao et al., 2025; Jiang et al., 2025). Existing work offers only partial, largely phenomenological explanations for each phenomenon in isolation: plateaus and the collapse of solution diversity are linked to entropy exhaustion and over-exploitation of a fixed reward signal (Cui et al., 2025; Hao et al., 2025; Jiang et al., 2025); the initial shortening of solutions is attributed to reward-driven pruning of redundant reasoning, whereas the subsequent lengthening has been ascribed either to a shift toward solving harder instances that demand longer chains of thought, or to the emergence of multi-pass, self-reflective reasoning (“aha”-like behavior) under RLVR(Fatemi et al., 2025; Wang et al., 2025); and catastrophic forgetting is analyzed as an objective-mismatch problem between RLVR and subsequent SFT or other post-training stages (Luo et al., 2025b; Ding & Wang, 2025). However, these accounts remain fragmented: we still lack a unifying framework that ties all four behaviors to a common underlying structure or dynamical mechanism of RLVR-trained policies, and explains why they so often arise together in practice.
@@
A growing body of work tries to explain these phenomena by modeling LLM’s reasoning process as an implicit reasoning graph—a network of semantic states and logical transitions (Wang et al., 2024; Dutta et al., 2024; Xiang et al., 2025; Cabannes et al., 2024; Minegishi et al., 2025). However, translating this idea into a fully microscopic analysis has proven extremely difficult: constructing the reasoning graph from a model’s high-dimensional latent space is a formidable open problem, preventing a direct investigation into the structural origins of the observed dynamics (See Appendix C for details).
@@
Here, we argue that the very universality of the above RLVR dynamics provides a path forward. In complex systems, collective emergent behaviors are largely independent of microscopic details and governed by the system’s large-scale organization (Anderson, 1972; Watts & Strogatz, 1998; Barabási & Albert, 1999). This insight motivates an approach rooted in the core logic of the Renormalization Group (RG) (Wilson & Kogut, 1974): instead of directly studying the system at the full microscopic level, one can model and approach the system at a coarse-grained level to explain its macroscopic behavior. Applying this strategy, we shift our focus from the full reasoning graph at the token level to its coarse-grained backbone at the semantic level, which emerges by abstracting away fine-grained skills encoded in local network patterns like motifs or community structures (Milo et al., 2002; Fortunato, 2010; Zhang & Moore, 2014). This leads us to our central hypothesis regarding the emergent structure that forms under RLVR training for general-purpose System-2 reasoning:
@@
Figure 1: CoNet Reproduces Core LLM Training Dynamics. The minimal CoNet model (a) reproduces the two core empirical signatures of RLVR training observed in the DeepScaleR-1.5B LLM (b). These signatures are: (i) a two-stage reward dynamic consisting of a fast-learning stage followed by a slow-learning phase (top panels), and (ii) a non-monotonic, V-shaped evolution of the correct response length (bottom panels). This striking correspondence validates CoNet as a valuable minimal model for our theoretical analysis. The inset in (a) schematically depicts the CoNet abstraction. Lines in (b) are smoothed with a 60-step moving average.
To investigate this hypothesis, we adopt a comparative approach connecting macroscopic phenomenology to graph topology. We first confirm that the DeepSeek-R1-Distill-Qwen-1.5B model (Luo et al., 2025a) exhibits the aforementioned RLVR signatures: a two-stage learning curve, a V-shaped response trajectory, catastrophic forgetting, and policy collapse. We then show that these complex behaviors are qualitatively reproduced by the Concept Network Model (CoNet) (Cai et al., 2025), a minimal framework abstracting away neural implementation details. This striking resemblance suggests that these phenomena are emergent collective behaviors rather than artifacts of specific neural architectures, allowing them to be studied via a coarse-grained proxy. Employing CoNet, we identify the emergence of a sparse concept web that orchestrates these dynamics. This phenomenological alignment supports the hypothesis that similar topological self-organization underlies LLM reasoning, enabling us to derive the following mechanisms from the principles of sparse graph growth:
First, the V-shaped response length implies a structural evolution from local “skill islands” to a global concept web. The initial decrease results from the parallel optimization of independent skills, where the policy rapidly discards inefficient exploratory detours to converge on short, optimal paths for simple tasks (Fig. 2). The subsequent rise, however, marks the onset of global integration, where a tree-like web emerges (Fig. 3). On such a sparse substrate, the average geodesic distance is inherently correlated with network size. Thus, as the web expands to bridge distant islands during the slow-learning phase, the policy is geometrically forced to traverse longer paths. The rising V-curve is therefore not a change in reasoning style, but the result of the concept web’s continuous growth.
Second, the sparse topology explains the model’s specific vulnerability to catastrophic forgetting (Fig. 4). Unlike dense networks, a sparse web relies on critical “trunk” edges to support vast downstream branches. We show that SFT-induced forgetting acts as a topological disconnection: gradient updates overwrite weights at these high-traffic bottlenecks, severing bridges to entire sub-trees. This offers a geometric rationale for the “easy-to-break, easy-to-fix” nature of RLVR models; since the disconnected downstream knowledge remains intact, a brief RLVR resumption leads to rapid recovery by merely “re-soldering” the broken trunks rather than relearning the web.
Third, we trace policy collapse to the dynamics of learning at the web’s frontier (the leaf nodes). Extending the critical learning theory of Cai et al. (2025), our experiments suggest that during the slow-learning phase, a significant portion of tasks sequentially undergo continuous-phase-transition-like learning dynamics (Fig. 5). For these tasks, the policy shifts from broad exploration to locking onto specific solution paths. While this collapse is necessary for mastering individual skills, standard RLVR accumulates these local entropy reductions across the leaf nodes. This aggregation ultimately results in a global loss of diversity, manifesting as the policy collapse observed in reasoning benchmarks (Fig. 6).
Finally, identifying a “maximally frustrated state” at the onset of slow learning—where competition between disconnected islands peaks—motivates our Annealed-RLVR algorithm. By introducing a targeted SFT “heating” step at this topological bottleneck, we temporarily relax competitive tension, helping the system navigate past local optima before cooling into a refined structure. Our experiments confirm that this theory-driven approach resolves the bottleneck and achieves superior reasoning performance on in-distribution and out-of-distribution benchmarks (including Minerva and AIME), surpassing standard RLVR baselines by engineering the topology of reasoning.
@@
Figure 2: Visualizing the Structural Evolution from “Skill Islands" to the Concept Web. These network snapshots provide a direct, qualitative narrative of the concept web’s formation, corresponding to the different phases of training. In each subfigure, green, red, and blue dots indicate the question, answer, and intermediate nodes of the CoNet, respectively. The color of each directed edge represents the transition probability from the head to the tail and is drawn according to the colorbar in the rightmost column. Here, only edges with transition probabilities greater than $0.95$ are shown. (a) In the early fast-learning phase (Step 20), the model discovers a few short, disjointed reasoning paths, representing the first nascent “skill islands”. (b) At the onset of slow learning (Step 50), the system has proliferated into a maximal collection of disconnected islands. (c) Deep in the slow-learning phase (Step 800), these previously separate islands have coalesced into a single, giant connected component, forming the unified and expansive concept web.
@@ 2 The V-Shaped Signature as Evidence for a Sparse Concept Web
The V-shaped evolution of the response length is a well-demonstrated, quantifiable hallmark of deep structural reorganization during RLVR training (He et al., 2025; Luo et al., 2025a). To anchor our investigation, we first reproduce this signature by applying RLVR to the DeepSeek-R1-Distill-Qwen-1.5B model following the DeepScaleR protocol (Luo et al., 2025a). As shown in Fig. 1, our experiment recovers the characteristic V-shaped curve (Fig. 1).
To dissect the mechanism underlying this signature—a task intractable within the LLM itself—we deploy an extended, multi-task variant of the CoNet as a “computational microscope” (See Appendix D for details). CoNet instantiates the implicit reasoning graph discussed in the Introduction as a concrete, tractable model: it maps the LLM “semantic states” to abstract nodes in a fixed random graph and LLM “logical transitions” to learnable, probabilistic edges. This minimal model reduces the learning process to a tractable graph-traversal problem and, remarkably, reproduces the V-shaped curve observed in our LLM experiment (Fig. 1), validating it as a faithful proxy for uncovering the underlying mechanism.
@@
Figure 3: A Sparse-Web Structure Necessitates Longer Reasoning Chains. This figure links the emergent sparse topology of the concept web to the observed increase in response length. (b, c) The color and marker scheme is identical to that used in Fig. 2, and only edges with transition probability greater than $0.95$ are retained. The largest connected component (a local region of which is shown here) remains sparse with average degree around 22 even as it grows from step $50$ to $800$. (a) Consequently, the distribution of solution path lengths, shown in both raw counts (histogram) and probability density (inset), shifts decisively to the right during this period, confirming that navigating the sparse backbone requires longer reasoning chains.
Armed with this confirmation, we now demonstrate that the V-shaped evolution of the response length arises from a two-stage process. The decreases of the response length in the initial fast-learning stage is dominated by local optimization. Given the random initialization of transition probabilities, shorter paths inherently have a higher probability of being sampled. Consequently, the model rapidly discovers short, efficient solutions to individual problems, giving rise to a proliferation of disconnected “skill islands” (as shown in Fig. 2(a)). This burst is quantified by the number of distinct solution clusters, which peaks right before the onset of slow learning (Fig. 2(b)). Stage one thus constitutes a parallel search for local efficiency that naturally shortens the average response length.
@@
The subsequent slow-learning phase, where response length steadily increases, marks a fundamental shift to global integration. The primary task is no longer discovering new islands but weaving them into a single, expansive concept web (see Fig. 2(c)), defined to be the largest connected component in which every edge has a transition probability above $0.95$. As the concept web relentlessly grows, a critical structural property emerges: the web remains persistently sparse, with its average degree stabilizing near two (as shown in Fig. 3(c)). This provides direct, quantitative validation for our central Hypothesis (Hypothesis 1).
@@
This emergent sparsity is the topological key to the V-shape’s rising slope. A network with such a low average degree is dominated by chain and tree-like structures, inherently lacking the shortcuts and redundant pathways found in denser graphs. Consequently, connecting two previously distant concepts requires traversing long, intermediate paths. The histograms in Fig. 3(a) provide clear evidence of this effect: As the sparse web expands, the distribution of solution lengths undergoes a decisive and sustained rightward shift, confirming that navigation on this sparse backbone necessitates longer reasoning chains. Thus, the V-shaped curve furnishes a robust macroscopic signature of an underlying topological shift: from parallel, local optimization on disconnected components to serial, global navigation across an expanding, sparse backbone.
@@
Yet can this simple, largely tree-like model capture the full complexity of LLM reasoning? After all, an LLM’s autoregressive memory enables sophisticated behaviors such as self-correction or reflection (e.g., “Wait, let me rethink that…”)—that would manifest as cycles in the reasoning graph. We argue they are fully compatible(Minegishi et al., 2025). These reasoning cycles are not the dense, ubiquitous shortcuts of a highly connected graph. Instead, they are likely the cognitive manifestation of sparse, local loops—occasional, high-cost revisions on a much larger, fundamentally chain-like scaffold. The existence of a limited number of such cycles does not significantly alter the global average degree of the web, which remains close to two. With the sparse-web model established, this link between a macroscopic dynamic and a microscopic topology forms the foundation for the analysis that follows.
@@
Figure 4: The Fragility of a Sparse Web: Catastrophic Forgetting and Fast Recovery. This figure demonstrates a key prediction of our sparse-web hypothesis: that the concept web is fragile, relying on critical bridge-like connections. (c-e) The color and shape conventions follow those of Fig. 2, and only a local view of the concept web is shown. Orange crosses on the edges indicate decreases in transition weights, with thicker crosses denoting more substantial decreases. The microscopic view in CoNet shows how a lightweight supervised fine-tuning (SFT) on a converged web (c) severs these critical bridges, causing the structure to fragment (d). Subsequent RLVR rapidly repairs these links (e). (a, b) This microscopic severing manifests as macroscopic catastrophic forgetting in both CoNet and the 1.5B LLM, where performance plummets upon initiating SFT. The subsequent fast recovery once RLVR resumes highlights the localized nature of the damage, underscoring the web’s structural fragility.